OpenStudy (maheshmeghwal9):

Q: - Equal masses of oxygen, hydrogen, & methane are taken in identical conditions. Find the ratio of the volumes of the gases.

OpenStudy (maheshmeghwal9):

@apoorvk @Callisto Please help:)

OpenStudy (apoorvk):

well, lets say 'x' grams of each are taken. Now am assuming ideal behavior by each gas, that is, equal moles of each gas occupies equal volume. (i am also assuming that's it's oxygen 'gas', i.e. O2 that is being taken, as well as H2) Now, Molar mass of O2 = 32 g/mol MM of H2 = 2 g/mol and Mm of CH4 = 16 g/mol So, in 'x' grams, there are x/32 moles of O2, x/2 moles of H2 and x/32 moles of CH4 present. You can find the ratio of moles, and from law of constant proportions, the ratio of volumes would be the same as the ratio of moles of each of compound!!

OpenStudy (maheshmeghwal9):

in CH4 isn't there x/16 moles??????????

OpenStudy (apoorvk):

oops yes, "x/16". Typo, sorry my bad.

OpenStudy (maheshmeghwal9):

k! so ratio is 16:1:8 ?

OpenStudy (apoorvk):

umm actually, no. you have \(\large \frac{x}{32}:\frac{x}{2}:\frac{x}{16}\) just multiply evry term by "32/x" and and you will get the required ratio. (I guess you got a bit confused up there).

OpenStudy (apoorvk):

*every

OpenStudy (maheshmeghwal9):

oops sorry ya i gt a bit confused there. thanx a lot I gt now so the ratio is 1:16:4

OpenStudy (maheshmeghwal9):

sorry 1:16:2

OpenStudy (maheshmeghwal9):

?

OpenStudy (apoorvk):

Yeah that's right-o!! :)

OpenStudy (maheshmeghwal9):

:)

OpenStudy (maheshmeghwal9):

thanx

OpenStudy (apoorvk):

No worries!!! Now do the Samba!!!! O.O

OpenStudy (maheshmeghwal9):

what is this "Samba"?

OpenStudy (maheshmeghwal9):

sorry my chat is poor:/

OpenStudy (maheshmeghwal9):

?

OpenStudy (apoorvk):

You dunno SAMBA??? It's just a dance form. Except that you will have to do it alone :P Just trolling mate hehe ;)

OpenStudy (maheshmeghwal9):

oh i see:) lol